OpenStudy (3psilon):

rotate the region bounded by y = 6 - 2x - x^2 and y = x+6 about the line y=3

OpenStudy (3psilon):

My answer differs from that of my teacher's. Can someone just show me the integral

zepdrix (zepdrix):

Can you provide the teacher's answer a sec? I wanna see if I made some terrible mistake before I post this integral :P

OpenStudy (3psilon):

ok ok

OpenStudy (3psilon):

Problem 21 part D

zepdrix (zepdrix):

Hmm this is setup really strangly... Clearly the parabola is our upper function. So why is the parabola being subtracted from the line... This is what I came up with, \[\large \pi \int\limits_{-3}^0 \left[(6-2x-x^2)-3\right]^2-\left[(x+6)-3\right]^2\;dx\] I should look over my work again though.. bit of a tricky problem. I may have made a mistake somewhere.

OpenStudy (3psilon):

But I thought the outer radius was at a height 3 plus the parabola function so why wouldn't it be 3+(6-2x-x^2)?

zepdrix (zepdrix):

So the outer radius \(\large R\) is going to be \(\large y_1-3\). Where \(\large y_1=6-2x-x^2\).